OpenStudy (angelwings996):

Consumer math help please: You invest $1,050.00, buying stock in a company at $10.00/share. The shares increase in value by 13% the first year and decrease 5% the second year. How much is the total value of your stocks in the second year?

OpenStudy (anonymous):

$1127.175

OpenStudy (angelwings996):

AlI I know to do is $1,050.00 divided by $10.00 which is 105

OpenStudy (anonymous):

@ayesdeeef is corret i got the same thing

OpenStudy (anonymous):

105 shares * 10.74 = 1,127.70

OpenStudy (anonymous):

Your welcome, don't forget to give a medal.

jimthompson5910 (jim_thompson5910):

"The shares increase in value by 13%" So focusing on one share, you go from $10 to ???

OpenStudy (angelwings996):

wait, do I do 10 * 13% ?

jimthompson5910 (jim_thompson5910):

so what is 13% of 10?

jimthompson5910 (jim_thompson5910):

you're on the right track in a way

OpenStudy (angelwings996):

1.3

jimthompson5910 (jim_thompson5910):

add that to 10 to get 11.3

jimthompson5910 (jim_thompson5910):

So a 13% increase makes you go from $10 to $11.30

jimthompson5910 (jim_thompson5910):

Now the stock's value drops by 5% So find 5% of 11.30 and subtract that from 11.30

OpenStudy (angelwings996):

okay I got 10.73

jimthompson5910 (jim_thompson5910):

That's the stock's price after year 2

jimthompson5910 (jim_thompson5910):

To find the total value, you multiply by 105 since you bought 105 shares

OpenStudy (angelwings996):

so it would be 1126.65

jimthompson5910 (jim_thompson5910):

wait a sec

jimthompson5910 (jim_thompson5910):

you got 10.735 right?

jimthompson5910 (jim_thompson5910):

not just 10.73

OpenStudy (angelwings996):

oh no I got just 10.73

jimthompson5910 (jim_thompson5910):

5% of 11.30 = 0.05*11.30 = 0.565 Subtract this from 11.30 to get 11.30 - 0.565 = 10.735

jimthompson5910 (jim_thompson5910):

Then you round that to the nearest penny to get 10.74

jimthompson5910 (jim_thompson5910):

Finally, multiply the individual share price by the number of shares: 10.74*105 = 1,127.70

OpenStudy (angelwings996):

Oh I get it. I rounded early so I did 11.30 - 0.57

jimthompson5910 (jim_thompson5910):

I see

jimthompson5910 (jim_thompson5910):

hmm now that I think about it, I'm not sure where rounding would be appropriate it is needed, but you may be right in saying you round early

jimthompson5910 (jim_thompson5910):

Let's go with your version since you're basically saying the stock loses 57 cents in value

OpenStudy (angelwings996):

Yeah cus I though you are supposed to round first then subtract.

jimthompson5910 (jim_thompson5910):

and often, rounding up is only done in favor of those who are in control

OpenStudy (angelwings996):

Yes, so would it be 1127.70 or 1126.65?

jimthompson5910 (jim_thompson5910):

but nvm about that, so yeah the answer is 1126.65

jimthompson5910 (jim_thompson5910):

I'd go with 1126.65 since rounding is usually done as soon as possible

OpenStudy (angelwings996):

Alright awesome thank you! Could you help me with just one more? I posted it earlier

jimthompson5910 (jim_thompson5910):

ok let me find it

OpenStudy (angelwings996):

Ill tag you in it. It was the one I posted before this one

jimthompson5910 (jim_thompson5910):

alright